Chemistry Lab
Interstitial water shipboard analysis on Leg 174A included refractometric analysis for salinity; titrations for pHm, alkalinity, and chloride; ion chromatography for sulfate, potassium, sodium, calcium, and magnesium; and colorimetric analyses for silica, phosphate, and ammonium. Atomic absorption spectrophotometry was used to quantify concentrations of Sr in pore waters.

Solid core samples were analyzed for inorganic and total carbon analysis (using the coulometer and the CNS). Based on their organic carbon content, some samples were selected and analyzed with the Rock-Eval. The system was used to determine S1, S2, and S3.

Gas chromatograph (GC)-3 was calibrated and used during this leg to provide real-time monitoring of the volatile hydrocarbons. Problems encountered on Leg 172 with the atomic absorption (AA) flame sputtering when using the air-acetylene/NO2 burner head were finally solved by lighting the flame without the aspirating water. Once the burner head heated enough to vaporize the aspirated water, it worked fine. Analysis of samples with low Sr levels are now possible. No or few problems were encountered with the remainder of the equipment in the lab.

Numerous suggestions and problems encountered previously with the coulometer program, the C1/C2 LabView application, the web, and JANUS have not been addressed.

Computer Services
A JANUS upgrade was installed in Halifax, and the two-way database updates were made. These procedures are time consuming and required the additional manpower that was available for this port call. Future port calls may require that an off-coming computer specialist remain with the ship for additional days to complete the tasks.

Some of the Business Objects User Reports run very slowly and need additional work to decrease their run time.

A Macintosh 840AV was used to broadcast the core log video display as a test. PC hardware has been requested; this is one less task residing on the VAX 3500. The VAXs are being phased out. The migration of MATMAN to FoxPro is in progress but not fast enough for the computer group, which does not want to expend time schooling new computer specialists in equipment that is being surplused. Also in a phase-out mode is the 4D database; Science Services' Equipment Status Report is now the only user.

Core Lab

Available time on this leg was used for some of the maintenance projects that develop over time. Some areas received touch-up painting, new cabinet baseboards, and cabinet repair. Loose faucets, noisy saw bearings, frozen linear bearings‹all received attention. H2S monitors, eye wash stations, and first aid stations were checked.


The low recovery experienced on this leg caused scientists to revise their sample requests because they were originally designed for high-recovery high-resolution studies. All requests were modified to correspond to the recovery at hand with the exception of the last site, Site 1073, where recovery approached the amount that was expected. This recovery left the plans for a high-resolution postcruise sampling party in limbo.

Time was spent learning why the Business Objects Users reports used by operations/engineering ran so slowly. A "custom view" preference used was found to be unnecessary. Removing this preference solved the problem.

Moving Science Services' leg data into Netscape queries to make them accessible on the ship web site has worked well and satisfied participating scientists. There is a possibility that Operations, which plans to stay with Business Objects Users reports, will still not have computer services support.

There was no problem with the JANUS curatorial applications.

Electronics Support
Daily support was provided to numerous small problems encountered when the systems were turned on, or as scientists became familiar with the equipment. Few major problems were encountered other than the ongoing attempt to get the second goniometer in the X-ray fluorescence (XRF) equipment to function properly. Routine maintenance of the copiers was accomplished. Assistance was provided to evaluate the WSTP Adara recording packages.

An ongoing project to remove TOTCO sensors and wiring from the rig was initiated.

Magnetics Lab
Leg 174A was a very quiet cruise for both paleomagnetics and the Core Lab, in general, because of the low core recovery. The Tensor tool was deployed on the last site.

Special Projects
€Tensor tool #2243 returned to the ship and deployed on Site 1073A.
€De-permed the mu-metal shields on the cryomag.
€Wound an external wire coil at the junction of the cryo dewar and degauss coils to minimize shield leakage.
€Backed up all the old VAX Paleomag account files and programs to the Uservol for future reference (legacy data migration).

Problems Encountered
€D-2000 discrete sample demagnetizer continued having problems detecting completion of the demag cycle. DTECH eventually sent the information needed to null the voltages in a zero field.

Microscope/Photography Labs
Science was supported without problems. Routine maintenance was performed.

Paleontology Lab
The lab accommodated five paleontologists this leg. Hydrofluoric acid was used by the palynologist without incident. An accumulation of sieves and centrifuge components were sorted and inventoried, and some were returned to shore for evaluation. Additional shelves were ordered for one cabinet in the paleo prep area to better organize the sieve collection. Both centrifuges were functioning after being repaired on the previous leg.

It is possible to overload the circuit that the lab dish washer uses, when ovens or hot plates are added. A heavy duty extension cord eliminated the problem, which is listed for attention.

Physical Properties
Physical properties measurements were collected with few problems with the computer controls or hardware. The LabView program controlling the MST worked flawlessly and was well received by its users. Uploads of data from the various instruments to JANUS were successful. Some familiar program problems remain but have been worked around.

Few problems occurred with the port call in Halifax or the shipments to and from that port. The authoriztion to ship hazardous materials through New York City could not be arranged, so logging sources and batteries remained on board for the next port.

Differential GPS service was contracted for this leg, primarily because of the shallow-water drilling. Support from the lab was to ensure ship position displays were available to DP and the drillers shack. This display was a comforting backup to the beacon display and could be used to maintain position if a beacon failed.

It was noted that the derrick shadowed the dGPS antennas when the ship passed through a heading of 200° to 125°, causing them to revert to GPS mode (and position scatter) and then reacquire the correction when the ship was better positioned.

Support was provided for the VSP experiments that were controlled and recorded by Schlumberger. Seismic Systems, Inc. of Houston provided the loaned GI gun sound source.

A new Ashtech GLONASS/GPS receiver was delivered by helicopter. It will provide better positioning capabilities world wide. See the Special Projects section under Miscellaneous.

The logging scientists utilized the underway plotter to good advantage, running recorded seismic profiles and reprocessing them as necessary.

X-ray Lab

Some X-ray diffraction (XRD) samples were prepared and analyzed for an assortment of clays and minerals. These samples were scanned with the Minolta color scanner to verify that dry photometric scans could provide more or better information than the wet scans done on the cores.

No samples were submitted for XRF analysis. Some standards were run primarily during troubleshooting efforts to get gonio 2 functioning correctly. The effort was frustrating as the problems were intermittent and sometimes seemingly unrelated. A repair call was scheduled for New York City.

The X-ray JANUS module was critiqued and many suggestions and problems were forwarded to Information Services for review.

To 174A Miscellaneous

To 174A Table of Contents

ODP Publications

ODP Homepage